# convSeq: Fast and Scalable Method for Detecting Patterns in Spike Data

### v1

@@ 1 Introduction
We introduce an efficient and scalable method for unsupervised detection of spatiotemporal patterns in neural activity based on optimizing a set of constrained 2D filters. Distinct from existing approaches (e.g. convNMF, seqNMF), we optimize the filters with backpropagation, which allows us to take advantage of popular automatic differentiation frameworks and GPU acceleration. To reduce the number of learnable parameters, we also propose an alternative formulation of our method, in which the filters themselves are parameterized as fixed-width truncated Gaussians. Our speed benchmarks show that the method, which we call convSeq, works significanly faster than existing pattern detection methods.
@@ 3.2 Visualization of structured spontaneous activity
Figure 1: Overview of our method (with $K=1$). Initially, no sequences are visible in the input data matrix $\mathbf{X}$ (A). After model fitting, the peaks in the convolution $\mathbf{\hat{x}}^{(k)}$ (D) of the optimized filter $\mathbf{W}^{(k)}$ (E) with the data matrix indicate the presence of patterns. Peaks in $\mathbf{\hat{x}}^{(k)}$ exceeding the significance threshold $\alpha$ (dotted line in D) indicate significant detections of the pattern occurrences. Then the rows of $\mathbf{W}^{(k)}$ are sorted according to the latency of maximum value (F). Finally, rearranging the rows of $\mathbf{X}$ with the obtained sorting indices exposes the sequences (B), which closely match the ground truth (C). The y-axis in all the panels except D corresponds to neuron IDs. Panles E and F are not to scale with Panels A-D.
@@ 4.3 Real data
In this experiment the position of the mouse was recorded simultaneously with the neural activity, and so we can verify the detected patterns against the ground truth – the ordering of neurons based on their known place fields. A neuron’s place fields are determined by measuring its activity across the environment: the more a neuron fires in a particular location, the more “preferred” that location is. As the animal goes though different locations on the track, neurons with similar place tuning are more likely to spike together, and this information can be used to rearrange the order of neurons to make place cell sequences clearly visible (Fig. 10B).
@@ 6 Notes on implementation, training and hyperparameter choice
Consider the case in which two patterns exist in the data, and one them is much stronger than the other. With $K=1$ the stronger of the two will be detected. With $K=2$, both patterns will be detected (i.e. the first, already detected, one and the second). Setting $K=3$ should result in still detecting the two patterns plus some spurious “pattern” by the third filter (whose convolution peaks should not reach statistical significance, because the cross-correlation term in Eq. 1 penalizes similar activations and, indirectly, filters that are tuned to the same pattern).
@@ Appendix A Hyperparameters
HP
Description
Value
$\beta_{\mathrm{xcor}}$
Diversity loss weight
0 if $K=1$ else 10.0 ($10^{5}$ in Fig. 20)
$\beta_{\mathrm{TV}}$
Total variation weight
15.2 in Fig. 1, 4.5 in Fig. 13, otherwise 100.0
$M$
Filter width (along the time dimension)
200 in Figs. 1 and 10, otherwise 100
lrate
Learning rate
0.1
$j$
Maximum cross-correlation distance
$M$
$\sigma$
Standard deviation of the filters’ Gaussians (2nd formulation)
16.0 (20.0 in Fig. 20

### v2

@@ 1 Introduction
We introduce an efficient and scalable method for unsupervised detection of spatiotemporal patterns in neural activity based on optimizing a set of constrained 2D filters. Distinct from existing approaches (e.g. convNMF, seqNMF), we optimize the filters with backpropagation, which allows us to take advantage of popular automatic differentiation frameworks and GPU acceleration. To reduce the number of learnable parameters, we also propose an alternative formulation of our method, in which the filters themselves are parameterized as fixed-width truncated Gaussians. Our speed benchmarks show that the method, which we call convSeq [^1] https://github.com/RomanKoshkin/conv-seq, works significanly faster than existing pattern detection methods.
@@ 3.2 Visualization of structured spontaneous activity
Figure 1: Overview of our method (with $K=1$). Initially, no sequences are visible in the input data matrix $\mathbf{X}$ (A). After model fitting, the convolution $\mathbf{\hat{x}}^{(k)}$ of the optimized filter $\mathbf{W}^{(k)}$ (E) with the data contains clearly visible peaks (D), suggesting the presence of repeating patterns. The peaks in $\mathbf{\hat{x}}^{(k)}$ exceeding the significance threshold $\alpha$ (dotted line in D) indicate significant detections. To make the patterns visible, the rows of the optimized filter $\mathbf{W}^{(k)}$ are sorted according to the latency of maximum value (F). Finally, rearranging the rows of $\mathbf{X}$ with the obtained sorting indices exposes the sequences (B), which closely match the ground truth (C). The y-axis in all the panels except D corresponds to neuron IDs. Panles E and F are not to scale with Panels A-D.
@@ 4.3 Real data
In this experiment the position of the mouse was recorded simultaneously with the neural activity, and so we can verify the detected patterns against the ground truth – the ordering of neurons based on their known place fields. A neuron’s place fields are determined by measuring its activity across the environment: the more a neuron fires in a particular location, the more “preferred” that location is. As the animal goes through different locations on the track, neurons with similar place tuning are more likely to spike together, and this information can be used to rearrange the order of neurons to make place cell sequences clearly visible (Fig. 10B).
@@ 6 Notes on implementation, training and hyperparameter choice
Consider the case in which two patterns exist in the data, and one of them is much stronger than the other. With $K=1$ the stronger of the two will be detected. With $K=2$, both patterns will be detected (i.e. the first, already detected, one and the second). Setting $K=3$ should result in still detecting the two patterns plus some spurious “pattern” by the third filter (whose convolution peaks should not reach statistical significance, because the cross-correlation term in Eq. 1 penalizes similar activations and, indirectly, filters that are tuned to the same pattern).
@@ 7 Conclusions and future directions
Acknowledgements
The first author thanks Ibrahim Alsolami for helpful discussions and acknowledges financial support from KAKENHI grant JP23KJ2131 and Google. We thank anonymous reviewers for their feedback and suggestions. The second author acknowledges support from KAKENHI grant JP23H05476.
@@ Appendix A Hyperparameters
HP Description Value
$\beta_{\mathrm{xcor}}$ Diversity loss weight 0 if $K=1$ else 10.0 ($10^{5}$ in Fig. 20)
$\beta_{\mathrm{TV}}$ Total variation weight 15.2 in Fig. 1, 4.5 in Fig. 13, otherwise 100.0
$M$ Filter width (along the time dimension) 200 in Figs. 1 and 10, otherwise 100
lrate Learning rate 0.1
$j$ Maximum cross-correlation distance $M$
$\sigma$ Standard deviation of the filters’ Gaussians (2nd formulation) 16.0 (20.0 in Fig. 20